complete the square, and find the roots of the quadratic equation.
x^2 + 16x = 0

Found 2 solutions by venugopalramana, stanbon:
Answer by venugopalramana(3286)

complete the square, and find the roots of the quadratic equation.
x2 + 16x = 0
X^2+2*X*8+8^2-8^2=0
(X+8)^2=8^2
X+8=+8...........OT..-8
X=0........OR........-16

Answer by stanbon(75887)

x^2 + 16x = 0
(x^2+16x+8^2) = 8^2
(x+8)^2 = 64
x+8 = +8 or -8
x=-16 or x=0
Cheers,
Stan H.